Write the equation of the parabola with vertex at (-2,3) and focus at (-2,-1).
Answer by lwsshak3(11628)

Write the equation of the parabola with vertex at (-2,3) and focus at (-2,-1).
Given parabola opens downward:
axis of symmetry: x=-2
Its basic form of equation: (x-h)^2=-4p(y-k), (h,k)=coordinates of vertex
p=4 (distance from vertex to focus on the axis of symmetry)
4p=16
equation: (x+2)^2=-16p(y-3)